The radio station Jack FM broadcast of 93.5 khz. Calculate its wavelength in nanometers (Hz = 1/s )

for the same radio station calculate the energy (E ) in kilojoules, kJ
Answer #1

Auswer Се ко», усе , а сі where, w is the premeny a 93.5 km hardest i s $93.5810HZ :ca velocity of light . Ac wavelength 2935Energy, Ech neplanens constant t 1626 X10M39.93:55*100 si E. J = 6.19810-29 = 6.19. 810-32KT is L a Energy of the radio stat